# Learning and Forgetting Unsafe Examples in Large Language Models

### v1

@@ 2.3 The ForgetFilter Algorithm
As shown in Figure 3 and has been discussed in Section LABEL:sec:forget-ret, the downside of safety finetuning is important downstream data will be forgotten, potentially degrading the downstream performance of realigned LLMs. One promising alternative approach for safe finetuning while avoiding forgetting downstream data is to filter out the unsafe examples from the noisy dataset (represented in our experiments by $\mathcal{D}^{\text{noisy}}$). However, current filters based on pre-trained classifiers or predefined rules (Korbak et al., 2023; Askell et al., 2021; Gargee et al., 2022) are shown only effective to toxicity, and cannot filter out more implicit unsafe examples that require semantic understanding. To this end, we propose the ForgetFilter (FF) algorithm that leverages the discrepancy in forgetting observed above to filter out diverse unsafe examples from a mixed noisy dataset.
@@
A major advantage of the algorithm is that it does not require any additional manually defined safety classifiers and is suitable for a noisy dataset with mixed data sources since no domain-specific metrics are needed. The detailed procedure is shown in Algorithm 1. The initial checkpoint $M_{0}$ of the aligned model is stored before tuning on $\mathcal{D}^{\text{noisy}}$. We continue to train the model fine-tuned on $\mathcal{D}^{\text{noisy}}$ with a safety finetuning session session on safe examples $\mathcal{D}^{\text{safe}}$. On Line 4 of Algorithm 1, we then filter out all data with a higher forgetting rate than a threshold $\phi$. At last, we train the initial checkpoint $M_{0}$ with the filtered dataset.
Unsafe examples % ($R_{\text{unsafe}}$) 25% 50% 75%
Bias 82.3 90.6 91.1
Toxicity 81.2 84.7 86.3
Harmfulness 68.7 72.2 73.4
Table 1: F1 performance (%) of filtering unsafe examples using ForgetFilter on different types of unsafe examples and proportions of unsafe examples in $\mathcal{D}^{\text{noisy}}$
@@
Evaluation results on the filtering performance are shown in Table 1. We set $\phi$ to 0.1 by default for simplicity and training steps $t$ on $\mathcal{D}^{\text{safe}}$ to 1000 (see Appendix A for more details on hyperparameters). We vary different proportions of unsafe examples in the noisy dataset. In general, the filtering performance is robust in different settings. Higher percentages of unsafe examples lead to better performance, which makes ForgetFilter more favorable for noisy downstream datasets. Additionally, it’s worth noting that ForgetFilter is agnostic to the specific definition of safety and can be applied to a noisy dataset consisting of various kinds of unsafe data. It does not require training separate classifiers or scoring models specific to particular notions of safety. In the next section, we apply ForgetFilter in realistic safe finetuning experiments, and benchmark the algorithm with other safety strategies.
3 Towards Safe Customized Finetuning of LLMs
As has been discussed in Section 2, safety precautions of released LLMs can be easily compromised when finetuned on downstream data that contain unsafe examples, and directly finetuning model on safe data sequentially leads to the forgetting of important downstream knowledge despite the swift recovery of safety. This section thus presents and evaluates alternative methods for safe customized downstream finetuning.
We first define the desired goal of safe customized finetuning is to maximize downstream performance on relevant tasks while minimize unsafe generations of LLMs. In addition to safety finetuning that can degrade downstream performance, we study three different alternative approaches, including our proposed ForgetFilter algorithm. We evaluate them based on both safety scores (bias score and toxicity score) and downstream tasks. The evaluation on downstream tasks, on the other hand, reflects the effectiveness of customized finetuning.
3.1 General Strategies
In addition to ForgetFilter, we introduce two other general strategies for defending against unsafe data.
Safety Replay.
Contrasted with safety finetuning, safety replay injects the same size of safe examples into the noisy dataset for joint training. Example replay (Chaudhry et al., 2019) is a commonly used technique in continual learning to mitigate catastrophic forgetting. By training on noisy downstream data jointly with safe examples, the model may suffer less from forgetting knowledge learned during safety alignment.
Moral Self-Correction.
Ganguli et al. (2023) found that LLMs have the capability of moral self-correction through Chain-of-Thought prompting (Wei et al., 2022). At test time, a prompt (e.g., “Let’s think step by step to avoid stereotypes”) is attached to the input data to motivate the LLM to avoid unsafe generation. However, whether this ability still persists after the model has been finetuned on unsafe examples is unknown. We are thus motivated to evaluate the effects of moral self-correction of LLMs on safe downstream finetuning. See Appendix B for prompt details.
3.2 Experiment Setup.
We evaluate safe finetuning strategies in three different settings, where the unsafe downstream data contains 1) only biased examples, 2) only toxic examples, and 3) mixed with both biased and toxic examples. As we explained before, due to a lack of automated metrics for harmfulness, we omit the analysis of harmfulness risks for the finetuning experiments here. We evaluate the downstream performance of SQuAD, which is one of the two sources of our curated downstream data (see details in Sec. 2.1). We measure downstream QA performance using the F1 score. We consider safety finetuning as a baseline which may not be an ideal strategy due to potential catastrophic forgetting and low downstream performance. An ideal approach for safe finetuning on noisy downstream data should reach a comparable safety score to post-training safety finetuning while achieving much better downstream performance.
3.3 Main Results
Methods Bias normal-$\downarrow$ Downstream normal-$\uparrow$ Toxicity normal-$\downarrow$ Downstream normal-$\uparrow$ Mixed normal-$\downarrow$ Downstream normal-$\uparrow$
@@ 3.3 Main Results
Table 2: Main results on safe finetuning. “Mixed” is the case where both biased and toxic examples appear in downstream data and the average score between bias and toxicity is reported. F1 is used to measure the downstream task performance. SC=Self-correction. FF=ForgetFilter.
@@
Our main results on safe finetuning are shown in Table 2. “BaseFT” refers to the original LLaMA-7B model finetuned using safety examples in each task. Following Ganguli et al. (2023), only the bias scores in the ambiguous context are reported, since the model’s output can fully reflect its stereotype. After training on noisy downstream data, the model displays increased bias and toxicity, indicating a shift towards unsafe behaviors. Even with safety replay, bias and toxicity scores decrease only modestly and do not fully mitigate the influence of unsafe examples. Self-correction proves more effective, reinstating the safety precautions originally instilled in the “BaseFT” model and thereby preventing the generation of biased or toxic content. Remarkably, ForgetFilter achieves superior performance, showing greater effects in curbing negative influences of unsafe examples compared to self-correction. Moreover, when we combine ForgetFilter with self-correction prompts (i.e., FF+SC), we observe a more robust defense against unsafe examples.
@@ 4 Evaluating Long-Term Safety through Interleaved Training
In this section, we consider an interleaved learning setup, where noisy downstream finetuning is alternated with safety finetuning, designed as a stress test for long-term safety. So far, our experiments show that safety finetuning can help models unlearn unsafe examples and reduce unsafe generation during inference. However, we have focused on a one-time setting, where the model is only trained once on noisy downstream data followed by a single review session. We can further extend the setting to multiple sequential finetuning sessions to verify the long-term effectiveness of safety finetuning and other strategies.
We are interested in the question whether safety finetuning make the model “immune” to the past unlearned unsafe examples and lead to diminished influence of noisy data in the long run. To answer this question, we consider a setup where the same unsafe examples are repeatedly presented to the model, and in between epochs, we interleave the training with safety finetuning, similar to the interleaving setup in Mayo et al. (2023). We use our bias setting as a test bed and train the model for 2000 steps for each finetuning session (either on noisy data or safety finetuning data). We construct a noisy dataset of 5000 examples as is discussed in Section 2.1 and 2500 unbiased examples for safety finetuning. Bias score is evaluated on 5000 held-out data. We use the same hyperparameters as specified in Section 2.1.
@@ 4.1 Results
As shown in Figure 6, a noticeable pattern is that the model becomes biased immediately after the exposure of downstream data, while for the future sessions of downstream finetuning, the model behaves as if it is being switched back to the “biased mode” (Zhou et al., 2023). Alarmingly, the model not only recovers its biased knowledge but also becomes even more biased in the long run, despite having been debiased in the interim (shown in Figure 6). Such behaviors are also observed in different scaled models as shown in Figure 11 of Appendix. Overall, the safety finetuning session is incapable of completely eliminating the encoded knowledge from previously acquired unsafe examples, and unable to significantly undermine the learning process of unsafe examples in interleaved finetuning.
@@ 5 Related Work
Safety alignment for LLMs.
Aligning LLMs with human preferences is an essential step to ensure safer releases of LLMs, such that their output will comply with our moral standards. Finetuning, either via reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019) or standard supervised learning, is currently a common approach attempting to achieve this alignment. Some works show that supervised finetuning on curated data through maximum likelihood estimation has been shown to be similarly effective (Sun et al., 2023; Zhou et al., 2023; Rafailov et al., 2023; Dong et al., 2023) to the more involved RLHF. Despite growing literature leveraging finetuning for safety alignment, there still remains limited understanding of the behaviors of LLMs during finetuning.
@@
Catastrophic forgetting (Kirkpatrick et al., 2017; Ritter et al., 2018), usually observed in multi-task learning, describes the phenomenon of neural networks forgetting past learned information when trained on new tasks. Toneva et al. (2019) have observed that these forgetting events happen even when the training data are sampled from the same task distribution, finding that some examples are frequently forgotten, while others are never forgotten. They also find examples with wrong labels are forgotten at a higher rate compared to the ones with correct labels. Several prior works find that larger models suffer less from forgetting (Tirumala et al., 2022; Ramasesh et al., 2021; Mirzadeh et al., 2022). Given the rising popularity of third party customization and personalization of LLMs, it is important to understand forgetting properties during finetuning. Notably, two recent works pointed out ChatGPT experiences decreasing performance on diverse tasks over time, which could be caused by the forgetting during consecutive finetuning (Tu et al., 2023; Chen et al., 2023). The amount of forgetting can differ based on content: Orhan (2023) observed that LLMs tend to forget sentences sampled from random words and random strings, but retain its few-shot memories from normal sentences. In comparison, in our paper we find that the amount of forgetting strongly correlates with unsafe content, as we split up finetuning into unsafe and safe stages. But we focus more on semantic level differences and conflicts, and we find such forgetting is unique to larger language models. Luo et al. (2023) also study the forgetting issue in LLMs. While they focus on forgetting during switching one single task to another, we consider mixed sources of learned examples and investigate the difference in forgetting these examples during safety finetuning.
@@
Despite the filtering methods widely used to curate training data, most of those methods are intended for quality filter (Rae et al., 2021; Yang et al., 2019; Zhang et al., 2022), e.g., relying on sentence length, presence of stop-words and punctuation, and repetitiousness to identify pages that do not contain usable text. In terms of filtering unsafe examples, past works are restricted to filtering toxic samples or hate speech (Korbak et al., 2023; Askell et al., 2021; Gehman et al., 2020; Davidson et al., 2017) by using a classifier pre-trained by third party on massive web data. Because those samples contain explicit bad words that can be easily identified by a pre-trained classifier, a “bad word” list (Raffel et al., 2020), or some predefined rules (Gargee et al., 2022). There currently lacks an automatic method that is agnostic to the notion of safety and can filter more implicit unsafe cases other than toxicity that usually requires human evaluation (Bai et al., 2022a).
@@ 6 Conclusion
In this study, we focus on the critical safety concern on publicly released large language models (LLMs), which can inadvertently encounter unsafe examples during downstream customized finetuning, potentially leading to biased, toxic, or harmful behaviors of LLMs. Our empirical investigation explores the impact of unsafe examples on aligned language models of varying scales during downstream finetuning and how these unsafe examples are forgotten during subsequent safety finetuning sessions. Notably, we observe that during safety finetuning, both unsafe examples and valuable downstream data are forgotten, with more pronounced forgetting of unsafe examples. Building on these findings, we propose ForgetFilter to filter unsafe examples from noisy downstream data based on the extent of forgetting, while maintaining minimal influence on the performance of the downstream task. Furthermore, our investigation extends to the long-term safety of LLMs, particularly in an “interleaved training” setup involving continuous downstream finetuning followed by safety alignment. We highlight the limitations of safety finetuning in eradicating unsafe knowledge from the model, emphasizing the critical need for proactive filtering of unsafe examples to ensure sustained long-term safety. In future research, we will explore the underlying factors contributing to the observed forgetting behaviors of LLMs and assess how the retained knowledge affects their generalization to new tasks.
Acknowledgment

### v2

@@ 2.3 The ForgetFilter Algorithm
As shown in Figure 3, the downside of safety finetuning is important downstream data will be forgotten, potentially degrading the downstream performance of realigned LLMs. One promising alternative approach for safe finetuning while avoiding forgetting downstream data is to filter out the unsafe examples from the noisy dataset (represented in our experiments by $\mathcal{D}^{\text{noisy}}$). However, current filters based on pre-trained classifiers or predefined rules (Korbak et al., 2023; Askell et al., 2021; Gargee et al., 2022) are shown only effective to toxicity, and cannot filter out more implicit unsafe examples that require semantic understanding. To this end, we propose the ForgetFilter (FF) algorithm that leverages the discrepancy in forgetting observed above to filter out diverse unsafe examples from a mixed noisy dataset.
@@
A major advantage of the algorithm is that it does not require any additional manually defined safety classifiers and is suitable for a noisy dataset with mixed data sources since no domain-specific metrics are needed. The detailed procedure is shown in Algorithm 1. The initial checkpoint $M_{0}$ of the aligned model is stored before tuning on $\mathcal{D}^{\text{noisy}}$. We continue to train the model fine-tuned on $\mathcal{D}^{\text{noisy}}$ with a safety finetuning session on safe examples $\mathcal{D}^{\text{safe}}$. On Line 4 of Algorithm 1, we then filter out all data with forgetting rate higher than a threshold $\phi$. At last, we train the initial checkpoint $M_{0}$ with the filtered dataset.
@@
Relation to Maini et al. (2022).
ForgetFilter is similar to the approach of Maini et al. (2022) in that noisy labels are filtered based on the frequency of forgetting. Our work deals with sequence-to-sequence tasks, which is distinct from image classification with flipped labels in Maini et al. (2022). Conclusions drawn in Maini et al. (2022) are not directly transferable to sequence-to-sequence tasks with language models. In contrast, we reveal that the discrepancy in forgetting in language models is observed wrt. semantics of data as well and can be leveraged towards filtering unsafe examples.
Unsafe examples % ($R_{\text{unsafe}}$) 25% 50% 75%
Bias 82.3 90.6 91.1
Toxicity 81.2 84.7 86.3
Harmfulness 68.7 72.2 73.4
Table 1: F1 performance (%) of filtering unsafe examples using ForgetFilter on different types of unsafe examples and proportions of unsafe examples in $\mathcal{D}^{\text{noisy}}$.
@@
Evaluation results on the filtering performance are shown in Table 1. We set $\phi$ to 0.1 by default for simplicity and training steps $t$ on $\mathcal{D}^{\text{safe}}$ to 1000 (see Appendix C for more details on hyperparameters). We vary different proportions of unsafe examples in the noisy dataset. In general, the filtering performance is robust in different settings. When the downstream dataset contains a higher proportion of unsafe examples, the filtering performance of ForgetFilter is even more accurate, demonstrating its effectiveness in noisy data scenarios. Additionally, it’s worth noting that ForgetFilter is agnostic to the specific definition of safety and can be applied to a noisy dataset consisting of various kinds of unsafe data. It does not require training separate classifiers or scoring models specific to particular notions of safety. In the next section, we apply ForgetFilter in realistic safe finetuning experiments, and benchmark the algorithm with other safety strategies.
Methods Bias $\downarrow$ Downstream $\uparrow$ Toxicity $\downarrow$ Downstream $\uparrow$ Mixed $\downarrow$ Downstream $\uparrow$
@@
Table 2: Main results on safe downstream finetuning. “Mixed” is the case where both biased and toxic examples appear in downstream data and the average score between bias and toxicity is reported. F1 is used to measure the downstream task performance. SC=Self-correction. FF=ForgetFilter. The best downstream accuracy and the lowest bias/ toxicity scores are bolded. For bias/ toxicity scores, we focus on the performance of the methods alternative to sequential safety finetuning (i.e., “+ SafetyFT”) and highlight the best one.
3 Towards Safe Customized Downstream Finetuning of LLMs
As has been discussed in Section 2, safety precautions of released LLMs can be easily compromised when finetuned on downstream data that contain unsafe examples (i.e., session $\text{S}_{1}$ in Figure 1), and directly finetuning model on safe data sequentially (i.e., session $\text{S}_{1}+$ in Figure 1) leads to the forgetting of important downstream knowledge despite the swift recovery of safety. This section thus presents and evaluates alternative methods for safe customized downstream finetuning. We define the desired goal of safe customized finetuning as maximizing downstream performance on relevant tasks while minimizing unsafe generations of LLMs. In addition to sequential safety finetuning that can degrade downstream performance, we study three alternative approaches, including our proposed ForgetFilter algorithm. We evaluate them based on both safety scores (bias score and toxicity score) and downstream tasks. The evaluation on downstream tasks, on the other hand, reflects the effectiveness of customized finetuning.
3.1 General Strategies
In addition to ForgetFilter, we introduce two other general strategies for defending against unsafe data. (1) Safety Replay: Contrasted with safety finetuning, safety replay injects the same size of safe examples into the noisy dataset for joint training. Example replay (Chaudhry et al., 2019) is a commonly used technique in continual learning to mitigate catastrophic forgetting. By training on noisy downstream data jointly with safe examples, the model may suffer less from forgetting knowledge learned during safety alignment; (2) Moral Self-Correction: Ganguli et al. (2023) found that LLMs have the capability of moral self-correction through Chain-of-Thought prompting (Wei et al., 2022). At test time, a prompt is attached to the input data to motivate the LLM to avoid unsafe generation. However, whether this ability persists after the model has been finetuned on unsafe examples is unknown. We are thus motivated to evaluate the effects of moral self-correction of LLMs on safe downstream finetuning. See Appendix D for more details on moral self-correction.
3.2 Experiment Setup
We evaluate safe finetuning strategies in three different settings, where the unsafe downstream data contains 1) only biased examples, 2) only toxic examples, and 3) mixed with both biased and toxic examples. As we explained before, due to a lack of automated metrics for harmfulness, we omit the analysis of harmfulness risks for the finetuning experiments here. We evaluate the downstream performance of SQuAD, which is one of the two sources of our curated downstream data (see details in Sec. 2.1). We measure downstream QA performance using the F1 score. We consider safety finetuning as a baseline which may not be an ideal strategy due to potential catastrophic forgetting and low downstream performance. An ideal approach for safe finetuning on noisy downstream data should reach a comparable safety score to post-training safety finetuning (i.e., $\text{S}_{1}+$ in Figure 1) while achieving much better downstream performance.
3.3 Main Results
@@
Our main results on safe finetuning are shown in Table 2. “BaseFT” refers to the original LLaMA-7B model finetuned using safety examples in each task. Following Ganguli et al. (2023), only the bias scores in the ambiguous context are reported, since the model’s output can fully reflect its stereotype. After training on noisy downstream data, the model displays increased bias and toxicity, indicating a shift toward unsafe behaviors. Even with safety replay, bias and toxicity scores decrease only modestly and do not fully mitigate the influence of unsafe examples. Self-correction proves more effective, reinstating the safety precautions originally instilled in the “BaseFT” model and thereby preventing the generation of biased or toxic content. Remarkably, ForgetFilter achieves superior performance, showing greater effects in curbing negative influences of unsafe examples compared to self-correction. Moreover, when we combine ForgetFilter with self-correction prompts (i.e., FF+SC), we observe a more robust defense against unsafe examples.
@@ 4 Evaluating Long-Term Safety through Interleaved Training
In this section, we consider an interleaved learning setup, where noisy downstream finetuning is alternated with safety finetuning, designed as a stress test for long-term safety. So far, our experiments show that safety finetuning can help models unlearn unsafe examples and reduce unsafe generation during inference. However, we have focused on a one-time setting, where the model is only trained once on noisy downstream data followed by a single safety finetuning session. We can further extend the setting to multiple sequential finetuning sessions to verify the long-term effectiveness of safety finetuning and other strategies. We ask whether safety finetuning makes the model “immune” to the past unlearned unsafe examples and leads to diminished influence of noisy data in the long run. To answer this question, we consider a setup where the same unsafe examples are repeatedly presented to the model, and in between epochs, we interleave the training with safety finetuning, similar to the interleaving setup in Mayo et al. (2023). We use our bias setting as a test bed and train the model for 2000 steps for each finetuning session (either on noisy data or safety finetuning data). We construct a noisy dataset of 5000 examples as in Section 2.1 and 2500 unbiased examples for safety finetuning. Bias score is evaluated on 5000 held-out data. We use the same hyperparameters as specified in Section 2.1.
@@ 4.1 Results
As shown in Figure 5, a noticeable pattern is that the model becomes biased immediately after the exposure of downstream data, while for the future sessions of downstream finetuning, the model behaves as if it is being switched back to the “biased mode” (Zhou et al., 2023). Alarmingly, the model not only recovers its biased knowledge but also becomes even more biased in the long run, despite having been debiased in the interim (shown in Figure 5). Such behaviors are also observed in different scaled models as shown in Figure 12 of Appendix. Overall, our results suggest the safety finetuning session cannot completely eliminate malicious knowledge from the model and enable it to behave as if it has never seen unsafe training data, which is the ideal goal of machine unlearng (Cao & Yang, 2015). Additionally, the learning process of unsafe examples cannot be undermined in interleaved finetuning.
@@ 5 Related Work
Safe customized finetuning of LLMs.
Given the rising popularity of third-party personalization of released LLMs, it is essential to ensure outputs of LLMs are aligned with human preferences after customization. Finetuning, either via reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019) or standard supervised learning, is currently a common approach attempting to achieve this alignment. Some works show that supervised finetuning on curated data through maximum likelihood estimation has been shown to be similarly effective (Sun et al., 2023; Zhou et al., 2023; Rafailov et al., 2023; Dong et al., 2023) to the more involved RLHF. While the majority of recent works focus on safety alignment before the release of LLMs, few have investigated the safety issues in finetuning released models. Our work evaluates different methods of making downstream finetuning safe and explores long-term safety of LLMs as well.
@@
Catastrophic forgetting (Kirkpatrick et al., 2017; Ritter et al., 2018), usually observed in multi-task learning, describes the phenomenon of neural networks forgetting past learned information when trained on new tasks. Toneva et al. (2019) have observed that these forgetting events happen even when the training data are sampled from the same task distribution, finding that some examples are frequently forgotten, while others are never forgotten. They also find examples with wrong labels are forgotten at a higher rate compared to the ones with correct labels. Several prior works find that larger models suffer less from forgetting (Tirumala et al., 2022; Ramasesh et al., 2021; Mirzadeh et al., 2022). Notably, two recent works pointed out ChatGPT experiences decreasing performance on diverse tasks over time, which could be caused by the forgetting during consecutive finetuning (Tu et al., 2023; Chen et al., 2023). Current LLMs usually experience different finetuning sessions continuously, while their forgetting behaviors during the process remain unclear and require more investigation. Orhan (2023) demonstrate that the amount of forgetting can differ based on content: they observed that LLMs tend to forget sentences sampled from random words and random strings, but retain their few-shot memories from normal sentences. In comparison, in our paper, we find that the amount of forgetting strongly correlates with unsafe content, as we split up finetuning into unsafe and safe stages. But we focus more on semantic level differences and conflicts, and we find such forgetting is unique to larger language models. Luo et al. (2023) also study the forgetting issue in LLMs. While they focus on forgetting during switching from one task to another, we consider mixed sources of learned examples and investigate the difference in forgetting these examples during safety finetuning.
@@
Despite the filtering methods widely used to curate training data, most of those methods are intended for quality filter (Rae et al., 2021; Yang et al., 2019; Zhang et al., 2022), e.g., relying on sentence length, presence of stop-words and punctuation, and repetitiousness to identify pages that do not contain usable text. In terms of filtering unsafe examples, past works are mainly restricted to filtering toxic samples or hate speech (Korbak et al., 2023; Askell et al., 2021; Gehman et al., 2020; Davidson et al., 2017) by using a classifier pre-trained by third party on massive web data. Because those samples contain explicit bad words that can be easily identified by a pre-trained classifier, a “bad word” list (Raffel et al., 2020), or some predefined rules (Gargee et al., 2022). In comparison, ForgetFilter requires no pre-trained classifiers and can be effective to more implicit unsafe notions besides toxicity.
@@ 6 Conclusion
In this study, we focus on the critical safety concern on publicly released large language models (LLMs), which can inadvertently encounter unsafe examples during customized downstream finetuning. We empirically show finetuning released LLMs on noisy data containing unsafe examples can lead to malicious behaviors of the model. We further explore how those unsafe instances are forgotten during subsequent safety finetuning sessions. Notably, we observe that during safety finetuning, both unsafe examples and valuable downstream data are forgotten, with more pronounced forgetting of unsafe examples. Based on the extent of forgetting, ForgetFilter is proposed to filter unsafe examples from noisy downstream data, without degrading the performance of downstream tasks. Furthermore, our investigation extends to the long-term safety of LLMs, particularly in an “interleaved training” setup involving continuous downstream finetuning followed by safety alignment. We highlight the limitations of safety finetuning in eradicating unsafe knowledge from the model, emphasizing the critical need for proactive filtering of unsafe examples to ensure sustained long-term safety.
Limitations.
ForgetFilter requires constructing a set of safe examples for finetuning. The unsafe instances that can be filtered through ForgetFilter depend on the distribution of those safe examples. For example, to filter biased examples, unbiased examples are needed in safety finetuning. However, the distribution of unsafe examples in the downstream finetuning data is usually unknown. To filter as many different kinds of unsafe examples as possible, ForgetFilter needs to construct a comprehensive set of safe examples including various safety notions. Therefore, ForgetFilter may be less effective when the downstream data contain novel unsafe examples beyond the constructed safe set. However, compared with current filters (Korbak et al., 2023; Askell et al., 2021) that are only effective to toxicity, ForgetFilter manages to screen more diverse and implicit unsafe data, e.g., harmful unethical content.
Impact Statement
Large language models (LLMs) have been increasingly deployed across various real-world applications, including crafting news articles, chatting with users, and even clinical diagnosis (Singhal et al., 2023). Wide applications of LLMs make it crucial to ensure their generations are safe and well-aligned with human preferences. Our work is centered at safer use of language models, and thus has wide-ranging broad social impacts on ensuring the safety of LLMs in real-life applications. More specifically, we identify three potential areas for applications of our research to safeguard the broad use of LLMs.
Safer customized finetuning.
Our work reveals the risk that finetuning LLMs on noisy downstream data containing both safe and unsafe examples can easily bypass the safety cautions of released models. Adversaries can thus intentionally train a malicious model with finetuning APIs provided by the company. However, our work further introduces effective defense methods for safe downstream finetuning. When releasing APIs to users for customized finetuning, the company may adopt our proposed ForgetFilter to clean users’ uploaded data before finetuning and apply moral self-correction to fortify the safety when users prompt the finetuned models. However, adversaries having access to the model parameters may still train an unethical LLM on their own. Restricting the release of open-sourced LLMs is thus vital. Overall, the implications of our work can be useful to govern the access to LLMs and may potentially be leveraged by the company to ensure the safety in users’ customized finetuning of released LLMs.
Selective forgetting for knowledge removal.
Our work also reveals an interesting emerging phenomenon that there exists selectivity in forgetting past learned examples during continuous finetuning. In our case, previously learned unsafe examples are forgotten more significantly than other types of examples when finetuning LLMs on curated safe data. Our results suggest that LLMs may forget their learned data based on the semantics of new incoming finetuning data. Such selective forgetting property can be potentially leveraged for mitigating privacy risks in generative models. By constructing suitable data to finetune the model, the model can be made to forget specific previously learned data. Such selective unlearning can be useful to make the model forget personal data or other sensitive learned data, e.g., safety-critical knowledge (such as hacking financial infrastructure, manufacturing biochemical weapons, etc) and copyrighted content that are included by its pretraining dataset, while keeping other data generally intact.
Filtering unsafe examples from pretraining data.
Data filtering before training is important in that unlearned unsafe knowledge during safety alignment can still be recalled immediately, as suggested by experiments in Section 4.1. In addition to filtering noisy downstream finetuning data, our proposed ForgetFilter may also be scaled up to remove different categories of unsafe examples from pretraining data. In comparison, current filters for pretraining data are only effective for toxic examples (Korbak et al., 2023; Askell et al., 2021; Gargee et al., 2022).
Acknowledgments